# MAFE: Multi-Agent Fair Environments for Decision-Making Systems

### v1

@@ 5.2 Reward Structure Customization
We design two types of rewards for agents: direct rewards and rate-based rewards. Direct rewards are explicit values, such as profits, that an agent aims to optimize. Rate-based rewards are expressed as ratios, such as the proportion of insured individuals to the total population, representing relative measures that agents aim to optimize. With this, we now provide the form of the reward summation in Problem 3.
@@
Let $K=j+l$, and define the reward components $[r_{1,t},...,r_{j+2l,t}]=c^{(R)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $r_{1,t},...,r_{j,t}$ are the direct rewards, $r_{j+1,t},...,r_{j+l,t}$ are numerators for rate-based rewards, and $r_{j+l+1,t},...,r_{j+2l,t}$ are denominators for the rate-based rewards at time $t$. Then, the final structure of the rewards summation in Equation 3 can be rewritten as the sum of its direct and rate-based constituents:
$\sum_{i=1}^{j}\alpha_{i}\left[\sum_{t=0}^{\infty}r_{i,t}\right]+\sum_{i=j+1}^{j+l}\alpha_{i}\left[\frac{\sum_{t=0}^{\infty}r_{i,t}}{\sum_{t=0}^{\infty}r_{i+l,t}}\right].$ (5)
5.3 Fairness Measure Structure Customization
@@
Given that the most common disparities in algorithmic fairness are rate-based, such as differences in insured rates across geographic regions in healthcare, we now describe how $F^{(m)}$ in Problem 3 is structured to measure these disparities when the number of groups is two or more.
@@
Two-group case. In the two-group case, the disparity between two groups is measured using the directly interpretable absolute difference in rates. Define the fairness components $[f_{1,t},...,f_{4M,t}]=c^{(F)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $f_{4m-3,t},...,f_{4m,t}$ represent the numerator and denominator for the rates of Groups 1 and 2 for the $m^{th}$ fairness measure. Then, the fairness violation is given by:
$F^{(m)}=-\bigg{|}\frac{\sum_{t=0}^{\infty}f_{4m-3,t}}{\sum_{t=0}^{\infty}f_{4m-2,t}}-\frac{\sum_{t=0}^{\infty}f_{4m-1,t}}{\sum_{t=0}^{\infty}f_{4m,t}}\bigg{|}$ (6)
@@
$D$-group case. When the number of groups, $D$, exceeds two, an absolute difference is inadequate for capturing disparities, as it fails to reflect the distribution of rates across multiple groups. To address this, we use standard deviation to quantify fairness disparities in the $D$-group case. Its simplicity provides an interpretable measure of how evenly rates are distributed among groups, making it particularly suitable for assessing fairness in multi-group settings. We define this measure as follows. Let the fairness components, $[f_{1,t},...,f_{2DM,t}]=c^{(F)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $f_{2D(m-1)+1,t},...,f_{2Dm,t}$, provide the numerator and denominator of each of $D$ groups for which we use for measuring the $m^{th}$ rate. Let $Y_{d}^{(m)}=\frac{\sum_{t=0}^{\infty}f_{2D(m-1)+d,t}}{\sum_{t=0}^{\infty}f_{2D(m-1)+d+1,t}}$ and $\mu^{(m)}=\frac{1}{D}\sum_{d=1}^{D}Y_{d}^{(m)}$. Then, the fairness measure is given by:
@@
6 Results and Analysis
(a) Loan Score Distributions
(b) Healthcare Score Distributions
(c) Education Score Distributions
Figure 2: Distribution plots that illustrate disparities in (a) the qualification score distributions of customers in the Loan Environment, (b) health risk score distributions among geographic sub-populations the Healthcare Environment, and (c) GPA score distributions of students in the Education Environment at the beginning of an episode.
(a) Debt Management
@@ 6 Results and Analysis
(b) Hospital Bed Availability
@@
(c) Insurance Availability
@@
(d) Public Health Investment
@@
(e) Tertiary Investment
@@
(g) Mentorship Programs
@@
(h) Employer Div. Incent.
@@
Figure 3: Plots illustrating the impact of various interventions in each environment, isolating their effects while holding other factors constant. (a) In the Loan MAFE, the effect of 20% debt relief on qualification scores for the full population. (b)-(d) In the Healthcare MAFE, the effects of providing hospital beds, universal health insurance, and unlimited public health investment on mortality rates. (e)-(g) In the Education MAFE, the effects of unlimited tertiary investment, full scholarships, and mentorship on graduation rates for the full population (e) and (f) and the disadvantaged population (g). (h) In the Education MAFE, the effect of unlimited diversity incentives for the Employer Agent on the average utility of workers from disadvantaged groups.
In Sections 6.1, 6.2, and 6.3, we demonstrate how MAFE actions address system disparities, verify agent learnability, and analyze agent action strategies. We introduce the Fair Multi-Agent Cross Entropy Method (F-MACEM) for optimizing these tasks, with details in Appendix B and additional results in Appendix C.
(a) Loan: Direct Rewards
(b) Loan: Fair Rewards
(c) Loan: Rate Rewards
(d) Healthcare: Direct Rewards
(e) Healthcare: Fair Rewards
(f) Healthcare: Rate Rewards
@@
Figure 4: Learning curves showing realized rewards obtained during training for models with different combinations of reward terms explicitly included in the F-MACEM’s objective function: “Direct"; “Direct + Fair"; or “Direct+Fair+Rate" in the objective.
6.1 Validating Interventions for Correcting Disparities
This section shows that actions shaped in our MAFEs can effectively mitigate disparities. Each MAFE is designed in a way that can incorporate structural biases, which may lead to disparate outcomes across demographic groups. The core attributes influencing outcomes vary by environment: qualification scores in the Loan MAFE, health risk scores in the Healthcare MAFE, and baseline GPA in the Education MAFE. These attributes reflect inherent biases across sensitive groups, calculated by regressing over dataset features used to construct each MAFE’s feature vectors. To enhance these biases for the purpose of supporting fairness research, we have resampled the original feature distributions, exacerbating disparities. Figure 2 illustrates these biased distributions at the start of each MAFE episode.
To assess whether agent actions can correct disparities, we conducted fixed intervention experiments, summarized in Figure 3. Using a fixed random seed, we compare the impact of specific interventions on environmental indicators with and without the intervention, repeating the process across five seeds. In the Loan MAFE, we examined debt management’s effect on qualification scores. In the Healthcare MAFE, we evaluated incidence and mortality rates under varying conditions like hospital bed availability, insurance coverage, and public health investments. In the Education MAFE, we analyzed the impact of investments, scholarships, mentorship programs, and diversity incentives on graduation rates and employer utility.
6.2 Compound Effects of Reward Terms
@@
In this section, we explore the cumulative impact of incorporating different terms into the F-MACEM’s objective function for each MAFE, specifically examining how various combinations of terms influence the observed outcomes for each individual term. We categorize these terms into three distinct groups, as outlined in Sections 5.2 and  5.3: direct rewards, fairness penalties, and rate-based rewards. To analyze their effects, we train the F-MACEM using three configurations of the objective function: (1) including only direct rewards, (2) including both direct rewards and fairness penalties, and (3) including direct rewards, fairness penalties, and rate-based rewards. For consistency, all elements in each objective function are uniformly weighted.
(a) Employer Agent
(b) Central Planner Agent
(General Intervention)
(c) University Budget Allocation Agent
Figure 5: Average actions taken by agents over training epochs in Education MAFE.
@@
The results of this analysis are presented in Figure 4. Each row corresponds to a different environment, while each column tracks the evolution of a specific reward category throughout training. Within each plot, the plotted curves differentiate the explicit reward terms included in the objective function. As expected, the red line—representing the objective function that explicitly incorporates all reward categories—shows steady improvement across all reward types during training. In contrast, configurations excluding certain terms often exhibit less consistent and volatile performance. For example, in the Education environment, the rate-based reward curve for the F-MACEM, trained solely with direct rewards, declines from its initial value during training and only approximately returns to its starting point by the final epoch on average. Similarly, in the Loan environment, excluding rate-based rewards causes the corresponding reward curve to plateau at a significantly lower value than observed in the fully-inclusive configuration. These patterns underscore the utility of integrating diverse reward terms to balance learning objectives effectively within each MAFE.
@@
This analysis also highlights environment-specific characteristics. Notably, the Healthcare MAFE shows smaller performance differences between training configurations compared to the Loan and Education MAFEs. While this might seem counterintuitive, it reflects the MAFE’s design: individuals transition between healthy, sick, and deceased states, with insurance profit as the primary reward. Insurers benefit most when the population maintains a high insured rate and remains healthy, minimizing claims. As a result, agents learn to balance interventions that optimize profitability and health outcomes. This alignment between agent objectives and system well-being offers a key insight: even when explicit stakeholder priorities diverge, overlapping indirect objectives can foster cooperative strategies that outperform narrow, self-serving approaches.
6.3 Policy Action Analysis
@@
In this section, we analyze the actions that the F-MACEM learns to produce over the training process when direct rewards, rate-based rewards, and fairness penalties receive uniform weighting in the objective function for the Education MAFE. We visualize how the Central Planner Agent distributes funds across different interventions, how the Employer Agent sets salaries, and how the university distributes the resources it receives for different interventions that improve student academic success.
Figures 5(a)-5(c) illustrate these agent actions. The Central Planner Agent primarily invests in tertiary resources and employer diversity incentives (Figure 5(b)), suggesting that tuition revenue adequately covers university operations. The University Budget Allocation Agent shifts its strategy the training process (Figure 5(c)). Initially, it allocates a significant portion of its budget to faculty salaries to ensure stability, but since faculty salaries are fixed, the agent refines its strategy by directing more resources to student-specific interventions, like scholarships and mentorship programs for underrepresented groups. This change helps reduce disparities in GPAs between majority and underrepresented students, improving overall educational and career outcomes.
Notably, Figure 5(a) shows a significant trend reversal in the Employer Agent’s salary-setting behavior midway through the training process. Initially, the Employer Agent decreases average salaries; however, this trend inverts as training progresses, leading to a steady increase in salaries. This shift results from multiple factors. First, the Central Planner Agent’s investment in diversity incentives directly boosts the salaries of underrepresented minority groups. Second, as the Central Planner Agent and University Budget Allocation Agent optimize their investments in tertiary resources and university student aid, overall student performance improves. These enhancements in educational outcomes translate to better career success, indirectly driving higher salaries.
Coordination among agents in each MAFE can create a positive feedback loop for improving system rewards, enabled by the flexible intervention structure our MAFEs offer. This structure facilitates the development of coordinated strategies that are more realistic and useful than the simplified abstractions used in single-action environments.
7 Conclusion and Discussion
In this work, we introduce the concept of Multi-Agent Fair Environments (MAFEs) as a framework for analyzing fairness in multi-agent systems. We provide a formal definition of algorithmic success within a MAFE, and develop three MAFEs modeling key social systems using a Python-based code implementation akin to popular reinforcement learning libraries like Gym, Gymnasium, and Petting Zoo. Through experimental analysis, we validate that our MAFEs can be used to analyze interventions that correct for system biases.
Fairness-aware algorithms require testing in environments that closely replicate real-world systems. While modeling human-centric systems involves some simplification, our MAFEs enhance the realism of decision-making in FairAI research. Acknowledging that domain experts may have varying perspectives on realism, our modular MAFEs offer flexible customization to meet diverse research needs. The models presented here represent one implementation, but our framework is adaptable and extensions will be analyzed in future work.
Disclaimer.
Appendix A Additional Related Work
Agent-based Social Simulations. Agent-based models (ABMs) have been employed to study various societal phenomena, such as the spread of misinformation in social networks, the propagation of epidemics, resource management, and economic systems (Perez & Dragicevic, 2009; Asgharpour et al., 2010; Giabbanelli et al., 2021; Benthall et al., 2021; Gausen et al., 2022). ABMs offer a bottom-up approach to understanding sociological phenomena, where the interactions between individual agents can lead to emergent behaviors (Elsenbroich & Polhill, 2023). Traditionally, such modeling has been conducted using surveys, network analysis, data mining, and game theory (Bonabeau, 2002). Recently, MARL has emerged as a powerful tool for analyzing complex group dynamics (Busoniu et al., 2008). However, the majority of existing MARL environments focus on specialized applications, such as games or autonomous navigation (Terry et al., 2021; Li et al., 2022) with limited relevance to fairness-oriented research. In contrast, our work analyzes fairness—an essential metric for assessing social and institutional interactions—in an MARL context.
Appendix B A Multi-agent Algorithm for Solving a MAFE
Algorithm 1 Fair Multi-Agent Cross Entropy Method (F-MACEM)
1:  repeat
3:     for episode $=1...$ number-of-episodes do
6:     end for
7:     Update $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}^{2}$ based on top $p\%$ of policies ranked by Equation 4.
8:  until Convergence
9:  Return $\boldsymbol{\theta}=\boldsymbol{\mu}$
In this section, we introduce the Fair Multi-agent Cross Entropy Method (F-MACEM), a simple yet effective algorithm for optimizing the objective function in Problem 3. The F-MACEM is an extension of the standard cross-entropy method (CEM), tailored to multi-agent systems with fairness considerations. This method is employed for performance analysis in Section 6.
Appendix C Additional Experiments
C.1 Policy Action Analysis
In this section, we provide the complete action analysis results from Section 6.3. Specifically, we present action analyses for each MAFE when direct, rate-based, and fairness violation terms are weighted uniformly in the objective function.
@@ C.1 Policy Action Analysis
For the Loan MAFE, we analyze the average admissions threshold set by the Admissions Agent, which determines the number of people approved for loans in an episode, and the debt management factor set by the Debt Management Agent, which helps the customer population avoid loan defaults. In the Healthcare MAFE, we examine how the Central Planner Agent allocates its budget across interventions and how the Insurance Agent sets premiums. For completeness, we restate the Education MAFE action analysis, focusing on how the Central Planner Agent distributes funds for interventions, how the Employer Agent sets salaries, and how the University Budget Allocation Agent allocates resources to improve student academic success.
@@
For the Loan MAFE, Figure 6(a) shows the average admission threshold over 40 training epochs. As training progresses, the agent learns to lower the threshold, effectively admitting nearly all applicants. This strategy increases the admission rate among the global population, thereby improving the rate-based reward. However, admitting more applicants without additional safeguards can increase default rates, risking the bank’s financial stability. To mitigate this issue, the Debt Management Agent can adjust the debt management factor to aid customers to avoid defaulting. As illustrated in Figure 6(b), this agent is able to strategically balance debt adjustment by setting these values neither too high to protect profits, nor too low to avoid widespread defaults. By targeting this aid, the agent ensures similar default rates across both groups, promoting fairness and financial stability.
@@
Figure 6(c) and 6(d) present the actions taken by various agents within the Healthcare MAFE. Specifically, Figure 6(c) highlights the premium-setting behavior of the Insurance Agent. During training, the agent learns to set premiums near the upper limit of \$1000. While this might initially seem challenging for affordability, Figure 6(d) illustrates a heatmap of the average percentage of the Central Planner Agent’s budget allocated to healthcare subsidies. The planner prioritizes two main areas: (1) subsidizing insurance premiums to reduce the effective cost for individuals and (2) investing in public health initiatives. These premium subsidies help maintain affordability for consumers, even with the higher nominal premiums. The largest share of the planner’s budget is allocated to public health investments, aimed at reducing the overall burden on the healthcare system by preventing illness. This approach focuses on improving baseline health outcomes across the population, complementing reactive measures like treatment subsidies by emphasizing preventive care strategies.
@@
Figure 6(e)-6(g) illustrate agent actions in the Education MAFE. The Central Planner Agent primarily invests in tertiary resources and employer diversity incentives, as shown in Figure 6(f), indicating that tuition revenue sufficiently covers university operations. The University Budget Allocation Agent demonstrates an evolving strategy, as shown in Figure 6(g). Early in the training process, the agent focuses a significant portion of its budget on faculty salaries to ensure financial stability and avoid potential disruptions. Yet, since faculty salaries in this MAFE are fixed, the agent recognizes that allocating too large a portion of its resources for them may not be the most efficient use of funds. As the agent refines its strategy, it adjusts its budget distribution, directing more resources toward student-specific interventions, such as scholarships for both majority and underrepresented student groups, as well as mentorship programs for underrepresented groups. This shift in allocation helps address disparities in cumulative GPAs between majority and underrepresented students, ultimately improving educational and career outcomes.
(a) Loan Frontier
(b) Healthcare Frontier
(c) Education Frontier
Figure 7: Pareto frontiers that demonstrate the reward-fairness tradeoff for the F-MACEM in the (a) Loan, (b) Healthcare, and (c) Education MAFEs.

### v2

@@ B.1 Reward Structure Customization
We design two types of rewards for agents: direct rewards and rate-based rewards. Direct rewards are explicit values, such as profits, that an agent aims to optimize. Rate-based rewards are expressed as ratios, such as the proportion of insured individuals to the total population, representing relative measures that agents aim to optimize. With this, we now provide the form of the reward summation in Problem 1.
@@
Let $K=j+l$, and define the reward components $r_{1,t},...,r_{j+2l,t}]=c^{(R)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $r_{1,t},...,r_{j,t}$ are the direct rewards, $r_{j+1,t},...,r_{j+l,t}$ are numerators for rate-based rewards, and $r_{j+l+1,t},...,r_{j+2l,t}$ are denominators for the rate-based rewards at time $t$. Then, the final structure of the rewards summation in Equation 1 can be rewritten as the sum of its direct and rate-based constituents:
$\sum_{i=1}^{j}\alpha_{i}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}\right]+\sum_{i=j+1}^{j+l}\alpha_{i}\left[\frac{\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}}{\sum_{t=0}^{\infty}\gamma^{t}r_{i+l,t}}\right].$ (2)
A concrete example of Equation (2) can be found in the healthcare MAFE description provided in Appendix G. For this environment, the direct reward term (the first summation in Equation (2)) corresponds to a single reward type: insurance profits. The rate-based reward term (the second summation) includes three types of rate-based measures: insured rates, negative incidence rates, and negative mortality rates.
B.2 Fairness Measure Structure Customization
@@
Given that the most common disparities in algorithmic fairness are rate-based, such as differences in insured rates across geographic regions in healthcare, we now describe how $F^{(m)}$ in Problem 1 is structured to measure these disparities when the number of groups is two or more.
@@
Two-group case. In the two-group case, the disparity between two groups is measured using the directly interpretable absolute difference in rates. Define the fairness components $f_{1,t},...,f_{4M,t}]=c^{(F)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $f_{4m-3,t},...,f_{4m,t}$ represent the numerator and denominator for the rates of Groups 1 and 2 for the $m^{th}$ fairness measure. Then, the fairness violation is given by:
$F^{(m)}=-\bigg|\frac{\sum_{t=0}^{\infty}\gamma^{t}f_{4m-3,t}}{\sum_{t=0}^{\infty}\gamma^{t}f_{4m-2,t}}-\frac{\sum_{t=0}^{\infty}\gamma^{t}f_{4m-1,t}}{\sum_{t=0}^{\infty}\gamma^{t}f_{4m,t}}\bigg|$ (3)
Both the Loan and Education MAFEs in Appendices F and H provide examples of the two-group sensitive attribute. In each environment, the sensitive attribute identifies whether a person belongs to a minority or majority demographic group. In the Loan MAFE disparities may arise between these groups with respect to key financial indicators, including admissions rates, average wait times, and and default rates. In the Education MAFE disparities may arise between these groups with respect to educational and career indicators, including university admissions rates, graduation rates, and average salaries.
@@
$D$-group case. When the number of groups, $D$, exceeds two, an absolute difference is inadequate for capturing disparities, as it fails to reflect the distribution of rates across multiple groups. To address this, we use standard deviation to quantify fairness disparities in the $D$-group case. Its simplicity provides an interpretable measure of how evenly rates are distributed among groups, making it particularly suitable for assessing fairness in multi-group settings. We define this measure as follows. Let the fairness components, $f_{1,t},...,f_{2DM,t}]=c^{(F)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})$, where $f_{2D(m-1)+1,t},...,f_{2Dm,t}$, provide the numerator and denominator of each of $D$ groups for which we use for measuring the $m^{th}$ rate. Let $Y_{d}^{(m)}=\frac{\sum_{t=0}^{\infty}\gamma^{t}f_{2D(m-1)+d,t}}{\sum_{t=0}^{\infty}\gamma^{t}f_{2D(m-1)+d+1,t}}$ and $\mu^{(m)}=\frac{1}{D}\sum_{d=1}^{D}Y_{d}^{(m)}$. Then, the fairness measure is given by:
@@
An example of a $D$-group sensitive attribute appears in the Healthcare MAFE in Appendix G. In this environment, geography serves as the sensitive attribute, and disparities may arise across four different geographic regions with respect to key health indicators, including mortality rates, incidence rates, and insured rates.
Appendix C A Multi-Agent Algorithm for Solving a MAFE
In this section, we introduce the Fair Multi-agent Cross Entropy Method (F-MACEM), a simple yet effective algorithm for optimizing the objective function in Problem 1. The F-MACEM is an extension of the standard cross-entropy method (CEM), tailored to multi-agent systems with fairness considerations. This method is employed for performance analysis in Section 6.
Appendix D Additional Experiments
D.1 Validating Interventions for Correcting Disparities (Unabridged)
This appendix section provides the unabridged version of Section 6.1 from the main body, including the MAFE-Health, MAFE-Loan, and MAFE-Edu results. Each MAFE is designed to incorporate structural biases, which may lead to disparate outcomes across demographic groups. The core attributes influencing outcomes vary by environment: health risk scores in the MAFE-Health, qualification scores in MAFE-Loan, and baseline GPA in the MAFE-Edu. These attributes reflect inherent biases across sensitive groups, calculated by regressing over dataset features used to construct each MAFE’s feature vectors. To enhance these biases for the purpose of supporting fairness research, we have resampled the original feature distributions, exacerbating disparities. Figure 5 illustrates these biased distributions at the start of each MAFE episode.
(a) MAFE-Health
(b) MAFE-Loan
(c) MAFE-Edu
Figure 5: Distribution plots that illustrate disparities in (a) health risk score distributions among geographic sub-populations in MAFE-Health, (b) the qualification score distributions of customers in MAFE-Loan, and (c) GPA score distributions of students in MAFE-Edu at the beginning of an episode.
Refer to caption (a) Hospital Bed Availability
@@ D.1 Validating Interventions for Correcting Disparities (Unabridged)
Refer to caption (b) Insurance Availability
@@
Refer to caption (c) Public Health Investment
@@
Refer to caption (d) Debt Management
@@
text
Refer to caption (e) Tertiary Investment
@@
text
text
Refer to caption (g) Mentorship Programs
@@
Refer to caption (h) Employer Div. Incent.
@@
Figure 6: Impact of various interventions in each environment, isolating their effects while holding other factors constant. (a)–(c) In MAFE-Health, the effects of providing hospital beds, universal health insurance, and unlimited public health investment on mortality rates. (d) In MAFE-Loan, the effect of 20% debt relief on qualification scores for the full population. (e)–(g) In MAFE-Edu, the effects of unlimited tertiary investment, full scholarships, and mentorship on graduation rates for the full population and the disadvantaged population. (h) In MAFE-Edu, the effect of unlimited diversity incentives for the Employer Agent on the average utility of workers from disadvantaged groups. Shaded regions provide standard deviations over random seeds.
To assess whether agent actions can correct disparities, we conducted fixed intervention experiments, summarized in Figure 6. Using a fixed random seed, we compare outcomes in environmental indicators with and without targeted interventions, repeating the process in five seeds. In MAFE-Health, we evaluated incidence and mortality rates under varying conditions such as hospital bed availability, insurance coverage, and public health investments. In MAFE-Loan, we examined debt management’s effect on qualification scores. In MAFE-Edu, we analyzed the impact of investments, scholarships, mentorship programs, and diversity incentives on graduation rates and employer utility.
D.2 Compound Effects of Reward Terms (Unabridged)
(a) Healthcare: Direct Rewards
(b) Healthcare: Fair Rewards
(c) Healthcare: Rate Rewards
(d) Loan: Direct Rewards
(e) Loan: Fair Rewards
(f) Loan: Rate Rewards
@@ D.2 Compound Effects of Reward Terms (Unabridged)
Figure 7: Learning curves showing realized rewards obtained during training for models with different combinations of reward terms explicitly included in the F-MACEM’s objective function: “Direct"; “Direct + Fair"; or “Direct+Fair+Rate" in the objective. Shaded regions provide standard deviations over random seeds.
@@
This appendix section provides the unabridged version of Section 6.2 from the main body, including the MAFE-Health, MAFE-Loan, and MAFE-Edu results. We particularly explore the cumulative impact of incorporating different terms into the F-MACEM’s objective function for each MAFE, specifically examining how various combinations of terms influence the observed outcomes for each individual term. We categorize these terms into three distinct groups, as outlined in Section 5: direct rewards, fairness penalties, and rate-based rewards. To analyze their effects, we train the F-MACEM using three configurations of the objective function: (1) including only direct rewards, (2) including both direct rewards and fairness penalties, and (3) including direct rewards, fairness penalties, and rate-based rewards. For consistency, all elements in each objective function are uniformly weighted.
@@
The results of this analysis are presented in Figure 7. Each row corresponds to a different environment, while each column tracks the evolution of a specific reward category throughout training. Within each plot, the plotted curves differentiate the explicit reward terms included in the objective function. As expected, the red line—representing the objective function that explicitly incorporates all reward categories—shows steady improvement across all reward types during training. In contrast, configurations excluding certain terms often exhibit less consistent and volatile performance. For example, in MAFE-Edu, the rate-based reward curve for the F-MACEM, trained solely with direct rewards, declines from its initial value during training and only approximately returns to its starting point by the final epoch on average. Similarly, in MAFE-Loan, excluding rate-based rewards causes the corresponding reward curve to plateau at a significantly lower value than observed in the fully-inclusive configuration. These patterns underscore the utility of integrating diverse reward terms to balance learning objectives effectively within each MAFE.
@@
This analysis also highlights environment-specific characteristics. Notably, MAFE-Health shows smaller performance differences between training configurations compared to MAFE-Loan and MAFE-Edu. While this might seem counterintuitive, it reflects the MAFE’s design: individuals transition between healthy, sick, and deceased states, with insurance profit as the primary reward. Insurers benefit most when the population maintains a high insured rate and remains healthy, minimizing claims. As a result, agents learn to balance interventions that optimize profitability and health outcomes. This alignment between agent objectives and system well-being offers a key insight: even when explicit stakeholder priorities diverge, overlapping indirect objectives can foster cooperative strategies that outperform narrow, self-serving approaches.
D.3 Assessing the Benefit of Multi-Agent Learning
In this section, we perform an experiment to assess the benefits of allowing multiple agents to learn dynamic policies, using the MAFE-Loan as a testbed. Specifically, we compare the performance of multi-agent learning, where all agents are allowed to learn optimal policies, against single-agent learning scenarios and a fixed policy baseline. The optimal policy, in this case, is defined as the one that maximizes the Loan MAFE’s objective function (as defined in equation 1), with uniform weighting applied to all terms in the objective.
D.4 Reward-Fairness Frontier in MAFEs
In this section, we analyze the F-MACEM algorithm’s performance in achieving fairness and accuracy, measured by the reward and fairness terms in Equation 1. Particularly, each reward and fairness violation is weighted uniformly, with $\alpha_{k}=\frac{\lambda}{K}$ for rewards and $\beta_{m}=\frac{1-\lambda}{M}$ for fairness violations. We then train the system using uniformly sampled values of $\lambda$ over the interval $[0,1]$ to analyze the trade-off between fairness and accuracy. To ensure uniform contribution from each component, we normalize all rewards and fairness violations to lie within the range $[0,1]$. The normalization factors for these results are provided in Table 8 of Appendix I.
(a) MAFE-Health Frontier
(b) MAFE-Loan Frontier
(c) MAFE-Edu Frontier
Figure 9: Pareto frontiers that demonstrate the reward-fairness tradeoff for the F-MACEM in the (a) MAFE-Health, (b) MAFE-Loan, and (c) MAFE-Health.
Figure 9 presents the resulting Pareto frontiers, which illustrate the trade-off between accuracy and fairness. Each point on the frontier represents the average performance of a model trained with the same objective function across three different training seeds to represent relative fairness values. Both fairness measures from Equation 3 (for Loan and MAFE-Edu) and Equation 4 (for MAFE-Health) produce negative values, which are plotted directly since they are compatible with maximization. In MAFE-Loan and MAFE-Edu, fairness is assessed using a binary sensitive attribute, with a higher value indicating greater fairness. In contrast, MAFE-Health evaluates fairness across four geographic regions, where a higher value also signifies greater fairness. In all plots, the highest fairness value corresponds to a value of 0.
D.5 Policy Action Analysis
@@
In this section, we analyze the actions that the F-MACEM learns to produce over the training process when direct rewards, rate-based rewards, and fairness penalties receive uniform weighting in the objective function for each MAFE.
@@
For the MAFE-Loan, we analyze the average admissions threshold set by the Admissions Agent, which determines the number of people approved for loans in an episode, and the debt management factor set by the Debt Management Agent, which helps the customer population avoid loan defaults. In the MAFE-Health, we examine how the Central Planner Agent allocates its budget across interventions and how the Insurance Agent sets premiums. For completeness, we restate the MAFE-Edu action analysis, focusing on how the Central Planner Agent distributes funds for interventions, how the Employer Agent sets salaries, and how the University Budget Allocation Agent allocates resources to improve student academic success.
@@
For the MAFE-Loan, Figure 10(a) shows the average admission threshold over 40 training epochs. As training progresses, the agent learns to lower the threshold, effectively admitting nearly all applicants. This strategy increases the admission rate among the global population, thereby improving the rate-based reward. However, admitting more applicants without additional safeguards can increase default rates, risking the bank’s financial stability. To mitigate this issue, the Debt Management Agent can adjust the debt management factor to aid customers to avoid defaulting. As illustrated in Figure 10(b), this agent is able to strategically balance debt adjustment by setting these values neither too high to protect profits, nor too low to avoid widespread defaults. By targeting this aid, the agent ensures similar default rates across both groups, promoting fairness and financial stability.
@@
Figure 10(c) and 10(d) present the actions taken by various agents within the MAFE-Health. Specifically, Figure 10(c) highlights the premium-setting behavior of the Insurance Agent. During training, the agent learns to set premiums near the upper limit of \$1000. While this might initially seem challenging for affordability, Figure 10(d) illustrates a heatmap of the average percentage of the Central Planner Agent’s budget allocated to healthcare subsidies. The planner prioritizes two main areas: (1) subsidizing insurance premiums to reduce the effective cost for individuals and (2) investing in public health initiatives. These premium subsidies help maintain affordability for consumers, even with the higher nominal premiums. The largest share of the planner’s budget is allocated to public health investments, aimed at reducing the overall burden on the healthcare system by preventing illness. This approach focuses on improving baseline health outcomes across the population, complementing reactive measures like treatment subsidies by emphasizing preventive care strategies.
@@
Figure 10(e)-10(g) illustrate agent actions in the MAFE-Edu. The Central Planner Agent primarily invests in tertiary resources and employer diversity incentives, as shown in Figure 10(f), indicating that tuition revenue sufficiently covers university operations. The University Budget Allocation Agent demonstrates an evolving strategy, as shown in Figure 10(g). Early in the training process, the agent focuses a significant portion of its budget on faculty salaries to ensure financial stability and avoid potential disruptions. Yet, since faculty salaries in this MAFE are fixed, the agent recognizes that allocating too large a portion of its resources for them may not be the most efficient use of funds. As the agent refines its strategy, it adjusts its budget distribution, directing more resources toward student-specific interventions, such as scholarships for both majority and underrepresented student groups, as well as mentorship programs for underrepresented groups. This shift in allocation helps address disparities in cumulative GPAs between majority and underrepresented students, ultimately improving educational and career outcomes.